# k-LLMmeans: Summaries as Centroids for Interpretable and Scalable LLM-Based Text Clustering

### v1

@@ Abstract
We introduce k-LLMmeans, a novel modification of the k-means clustering algorithm that utilizes LLMs to generate textual summaries as cluster centroids, thereby capturing contextual and semantic nuances often lost when relying on purely numerical means of document embeddings. This modification preserves the properties of k-means while offering greater interpretability: the cluster centroid is represented by an LLM-generated summary, whose embedding guides cluster assignments. We also propose a mini-batch variant, enabling efficient online clustering for streaming text data and providing real-time interpretability of evolving cluster centroids. Through extensive simulations, we show that our methods outperform vanilla k-means on multiple metrics while incurring only modest LLM usage that does not scale with dataset size. Finally, We present a case study showcasing the interpretability of evolving cluster centroids in sequential text streams. As part of our evaluation, we compile a new dataset from StackExchange, offering a benchmark for text-stream clustering.
k-LLMmeans: Summaries as Centroids for Interpretable and Scalable LLM-Based Text Clustering
Jairo Diaz-Rodriguez Department of Mathematics and Statistics York University Toronto, Ontario M3J 1P3 jdiazrod@yorku.ca
@@ 1 Introduction
Text clustering is a fundamental task in natural language processing (NLP), widely applied in document organization, topic modeling, and information retrieval (Schütze et al., 2008; Steinbach, 2000). A common approach involves generating text embeddings (Devlin, 2018; Sanh, 2019; Mikolov, 2013; Pennington et al., 2014; Brown et al., 2020) for each document, which are then clustered using traditional algorithms (Petukhova et al., 2025). Among these, k-means (MacQueen, 1967) is one of the most widely used, iteratively refining cluster centroids based on the mean of assigned points. Several alternative methods exist for computing centroids (Jain and Dubes, 1988; Bradley et al., 1996; Kaufman and Rousseeuw, 2008), yet they often rely on purely numerical calculations. While effective, such approaches risk losing crucial contextual and semantic information, as averaging high-dimensional embeddings can obscure nuanced textual relationships (Reimers and Gurevych, 2019).
@@
In this work, we propose a modification to the k-means algorithm by leveraging large language models (LLMs) to dynamically generate cluster centroids. We call this the k-LLMmeans. Instead of always computing centroids as the mean of embedding vectors, in spaced iterations we generate centroids based on textual summaries of the cluster contents. At each of these iterations, the centroid is represented as the embedding of the LLM-generated summary of the texts belonging to the cluster. Building on Jia and Diaz-Rodriguez (2025), the LLM-based centroids preserve key contextual and semantic aspects of the main documents in the cluster while filtering out secondary or lower-priority documents. This summary-based centroid provides a more interpretable and contextually relevant representation, mitigating the loss of meaning inherent in numerical averaging.
One key advantage of our approach is the interpretability of the clustering process. Traditional k-means lacks explicit explanations for how clusters evolve over iterations, whereas our method provides textual summaries that offer insights into the semantic shifts of cluster centroids. This can be practically beneficial for debugging and validating the expected behavior of the algorithm. Moreover, this transparency in current status of the clustering becomes a key advantage on sequential clustering where clusters might evolve on time and interpretability becomes important. In this area we provide a modified version of the mini-batch k-Means (Sculley, 2010) by utilizing our k-LLMmeans sequentially.
Several studies have explored unsupervised text clustering using LLMs (Zhang et al., 2023; Feng et al., 2024; De Raedt et al., 2023; Viswanathan et al., 2024; Shi and Sakai, 2023; Tarekegn et al., 2024; Nakshatri et al., 2023), demonstrating state-of-the-art performance across various datasets and benchmarks. These approaches consistently outperform traditional clustering algorithms such as k-means. However, a key limitation is their complexity and their reliance on fine-tuning or iterative querying of LLMs that scales with the dataset. While this improves clustering quality, it introduces instability, might requires extensive parameter/prompt tuning, and limits scalability for big data.
@@
Our approach is not designed to surpass complex state-of-the-art LLM-based text clustering methods but rather to provide a scalable and transparent LLM-enhanced modification to the well-established k-means algorithm, maintaining its key behavior. Fundamentally, a centroid is a numerical abstraction that represents a cluster, and we posit that an LLM-generated textual summary can serve an analogous role. Crucially, our approach preserves most of the mathematical properties of k-means, ensuring that core theoretical guarantees—such as convergence behavior, cluster compactness, and complexity— remain intact. Unlike alternative strategies that introduce complex pre-processing or post-processing steps, often altering the underlying mathematical framework, our method seamlessly integrates LLMs while maintaining k-means’ well-defined optimization landscape. This balance allows for enhanced interpretability and adaptability while preserving the efficiency that have made k-means a cornerstone of clustering algorithms.
@@
  • We introduce k-LLMmeans (Section 4), that leverages LLMs to enhance centroid estimation in k-means for text clustering.
  • We propose mini-batch k-LLMmeans (Section 5), designed for sequential, scalable, and interpretable text-stream clustering.
  • Through extensive simulations (Section 6), we demonstrate that both methods outperform k-means while maintaining low LLM usage that does not scale with the dataset size.
  • We present a case study (Section 7) demonstrating the interpretability of evolving cluster centroids in sequential text streams.
As a by-product of our experiments, we compile a dataset from StackExchange (2024) suitable for benchmarking text-streaming analysis methods (Section 6.2).
2 Related work
Clustering techniques are central to natural language processing and machine learning. Hierarchical methods (Johnson, 1967; Blashfield and Aldenderfer, 1978) build tree-structured representations of nested document relationships. Density-based approaches like DBSCAN (Ester et al., 1996) and graph-based methods detect clusters of arbitrary shapes, while spectral clustering (Ng et al., 2001) leverages eigen-decomposition to uncover complex structures. Model-based techniques—including Gaussian mixture models (Dempster et al., 1977) and recent neural network frameworks (Zhou et al., 2019; Huang et al., 2014; Yang et al., 2016; Zhang et al., 2021; Xie et al., 2016)—provide probabilistic clustering formulations. Additionally, topic modeling methods, from probabilistic latent semantic analysis (Hofmann, 2001) to latent Dirichlet allocation (Blei et al., 2003), capture word co-occurrence patterns and latent topics. However, these approaches diverge from our objective of enhancing k-means with LLMs, and are therefore not directly comparable.
Recent studies have integrated LLMs into clustering pipelines to reduce expert supervision and enhance performance. For instance, Viswanathan et al. (2024) employ LLMs to augment document representations, generate pseudo pairwise constraints, and post-correct low-confidence assignments for query-efficient, few-shot semi-supervised clustering. Similarly, Zhang et al. (2023) propose ClusterLLM, which uses instruction-tuned LLMs via interactive triplet and pairwise feedback to cost-effectively refine clustering granularity. Complementary approaches (Tipirneni et al., 2024; Petukhova et al., 2025) show that context-derived representations capture subtle semantic nuances beyond traditional embeddings. Additionally, Wang et al. (2023) introduce a goal-driven, explainable clustering method that employs natural language descriptions to clarify cluster boundaries, while De Raedt et al. (2023) present IDAS for intent discovery using abstractive summarization. Moreover, Feng et al. (2024) propose an iterative refinement mechanism that forms super-points to mitigate outliers and reassign ambiguous edge points, resulting in clusters with higher coherence and robustness. In contrast, our approach directly enhances the core k-means algorithm in an LLM-scalable manner.
@@ 2 Related work
             // Initialize centroids using k-means++
             // This step can be omitted if initial centroids are provided
@@
             // Update centroids with LLM every $l$ iterations
@@
             // Update centroids using standard averaging
@@
3 Preliminaries: k-Means for text clustering
We can start formalizing the text clustering problem. Given a corpus of $n$ text documents $D=\{d_{1},\cdots d_{n}\}$. Each document $d_{i}$ is represented as a $d$-dimensional embedding vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$ such that:
$\mathbf{x}_{i}=\text{Embedding}(d_{i}).$
We assume that all embeddings are normalized such that $\|\mathbf{x}_{i}\|=1$, equivalently using the euclidian distance or cosine similarity as the distance metric between embeddings. The goal of K-Means clustering is to partition these $n$ document embeddings into $k$ clusters, minimizing the intra-cluster variance. Formally, we define the clustering objective as:
$\arg\min_{C_{1},C_{2},\dots,C_{k}}\sum_{j=1}^{k}\sum_{i\in[C_{j}]}\|\mathbf{x}_{i}-\boldsymbol{\mu}_{j}\|^{2},$ (1)
$\boldsymbol{\mu}_{j}=\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}.$ (2)
The k-means algorithm assigns each document embedding $\mathbf{x}_{i}$ to the closest centroid based on the smallest distance and updates centroids accordingly until convergence after $T$ iterations. The primary objective of the k-means algorithm is to iteratively adjust cluster centroids to minimize the within-cluster variance, effectively guiding the algorithm toward an optimal set of centroids. However, due to its sensitivity to initialization and the non-convex nature of its objective function, k-means does not guarantee convergence to the global optimum and can instead become trapped in local optima (MacQueen, 1967; Lloyd, 1982). Various strategies, such as k-means++ initialization and multiple restarts, have been proposed to mitigate these issues and improve the likelihood of achieving better clustering results (Arthur and Vassilvitskii, 2006).
4 k-LLMmeans
@@ 3 Preliminaries: k-Means for text clustering
where $p_{j}=\text{Prompt}\left(I,\{d_{z_{i}}|z_{i}\sim[C_{j}]\}_{i=1}^{m_{j}}\right)$, and $m_{j}=\min(m,|C_{j}|)$. Here, $z_{i}\sim[C_{j}]$ denotes a sampled index of the embeddings assigned to cluster $C_{j}$ (without repetitions) and $m$ is a parameter that represents the maximum number of sampled indices used to compute the cluster centroid $\boldsymbol{\mu}_{j}$. In simple terms, we update a cluster’s centroid by using the embedding of the response generated by an LLM when queried with a prompt containing a summarization instruction $I$ and a representative sample of documents from the cluster. Rather than providing all documents within the cluster as input, the LLM processes a representative sample as a context prompt. While incorporating the entire cluster is theoretically possible, it poses practical challenges due to prompt length limitations. Therefore, we propose selecting the sample cluster documents using a k-means++ sampling of the cluster embeddings. Our experiments demonstrate that this sampling process facilitates a more effective synthesis of the cluster’s content, leading to improved summaries and, consequently, more refined centroid updates. The instruction $I$ varies depending on the clustering task, but standard summarization prompts are generally sufficient.
@@
This novel approach can help mitigate the tendency of k-means to get stuck in local optima by dynamically adjusting centroids using semantic insights from the data. Unlike standard k-means, which relies solely on Euclidean updates, an LLM can refine centroids based on contextual meaning and high-dimensional representations, allowing for better adaptation to complex structures in text data. By periodically re-centering clusters with LLM-informed embeddings, the algorithm can escape poor local minima and achieve more coherent and semantically meaningful clusters, even when the initial k-means++ seeding is suboptimal. Apart from this modification, our algorithm adheres to the core principles of k-means, preserving its well-established properties and ensuring practical robustness. We demonstrate in Section 6.4 that k-LLMmeans consistently outperforms k-means across extensive simulations.
4.1 Scalability and transparency
Our approach offers two key advantages over more complex LLM-based clustering methods: scalability and transparency. Unlike most state-of-the-art methods, whose LLM usage complexity grows with sample size (Feng et al., 2024), our algorithm depends only on $k$ and $l$, with even small $l$ values yielding performance gains over k-means. Additionally, our method enhances interpretability by producing LLM-updated centroids that meaningfully represent clusters, allowing practitioners to track and validate the algorithm evolution without requiring post-processing. While similar insights can be extracted indirectly in other clustering algorithms, our approach integrates this interpretability directly into the clustering process, improving both usability and practical applicability (See Figure 1 for an example of such evolution of summaries).
@@ 5 Mini-batch k-LLMmeans
input : $\ \{D_{1},\cdots,D_{b}\},k,I,m,l,T$
// $b$ batches of documents
@@
       // Update centroids proportional to current cluster sizes and batch cluster sizes
@@
Mini-batch k-means (Sculley, 2010) is an efficient strategy for large-scale text clustering that processes small, randomly sampled mini-batches instead of the full dataset. This approach substantially reduces memory usage and computational cost, making it well suited for continuously generated text streams—such as those from social media, news, or customer feedback—where data must be clustered incrementally without full dataset access. Mini-batch k-means exhibits convergence properties comparable to standard k-means while offering superior scalability.
@@
Although numerous streaming clustering methods that do not rely on LLMs have been studied (Silva et al., 2013; Aggarwal, 2018; Ribeiro et al., 2017; Aggarwal et al., 2003; Ackermann et al., 2012; Ordonez, 2003), only a few have incorporated LLMs (Tarekegn et al., 2024; Nakshatri et al., 2023). Moreover, existing offline LLM-based clustering approaches face scalability issues, highlighting the need for scalable LLM-driven clustering in an online setting. To address this gap, we introduce mini-batch k-LLMmeans, which directly extends mini-batch k-means by incorporating minimal LLM usage during centroid updates. Algorithm 2 details how mini-batch k-LLMmeans sequentially receives $b$ batches of documents $D_{1},\dots D_{b}$ where each batch contains a set of documents (these batches can either be random samples from a large corpus or represent sequential data). It processes each batch sequentially with k-LLMmeans and updates centroids incrementally using a weighted rule like mini-batch k-means. Our mini-batch k-LLMmeans algorithm preserves the desirable properties of mini-batch k-means, with low memory and LLM usage. Section  6 shows that it also outperforms in simulations.
6 Experiments
Dataset/Method Bank77 CLINC GoEmo Massive (I) Massive (D)
NMI dist NMI dist NMI dist NMI dist NMI dist
distilbert k-means 64.4 0.335 77.2 0.34 18.3 0.364 58.1 0.366 45.1 0.309
k-medoids 56.0 0.591 66.8 0.69 14.0 0.769 42.6 0.847 27.2 0.81
k-LLMmeans-1 64.8 0.351 77.4 0.352 18.3 0.358 58.6 0.362 46.0 0.294
k-LLMmeans-5 64.8 0.358 77.7 0.356 18.3 0.354 59.0 0.363 45.6 0.306
k-LLMmeans-FS1 64.8 0.352 78.1 0.346 18.9 0.357 58.8 0.365 45.7 0.299
k-LLMmeans-FS5 64.9 0.363 78.7 0.343 18.8 0.351 59.0 0.362 46.4 0.295
openai k-means 83.0 0.225 92.0 0.2 20.5 0.287 72.4 0.32 67.9 0.246
k-medoids 69.5 0.639 77.7 0.694 15.9 0.852 52.4 0.86 38.8 0.797
k-LLMmeans-1 83.6 0.221 92.5 0.194 22.3 0.278 73.0 0.314 69.6 0.232
k-LLMmeans-5 83.6 0.226 92.8 0.186 22.1 0.275 73.5 0.307 69.5 0.238
k-LLMmeans-FS1 83.8 0.219 92.8 0.186 21.9 0.283 73.6 0.314 69.4 0.233
k-LLMmeans-FS5 84.1 0.22 93.1 0.179 22.3 0.278 73.9 0.302 70.6 0.227
e5-large k-means 76.7 0.142 90.8 0.131 22.8 0.176 70.9 0.175 63.7 0.138
k-medoids 64.5 0.331 73.4 0.371 15.7 0.447 50.3 0.437 36.3 0.405
k-LLMmeans-1 78.6 0.138 91.6 0.127 23.4 0.176 72.0 0.17 64.8 0.134
k-LLMmeans-5 79.0 0.136 92.0 0.127 23.7 0.174 72.1 0.166 66.0 0.131
k-LLMmeans-FS1 78.8 0.136 91.7 0.127 23.7 0.174 71.6 0.173 64.9 0.136
k-LLMmeans-FS5 79.5 0.134 92.5 0.119 24.3 0.168 72.4 0.166 65.9 0.133
sbert k-means 80.9 0.255 91.0 0.215 13.3 0.355 70.7 0.344 64.6 0.271
k-medoids 69.0 0.64 78.5 0.68 12.0 0.899 54.6 0.817 44.6 0.8
k-LLMmeans-1 81.7 0.253 91.6 0.215 13.6 0.351 71.2 0.334 65.6 0.248
k-LLMmeans-5 81.8 0.253 91.9 0.21 13.7 0.348 71.5 0.325 65.2 0.25
k-LLMmeans-FS1 82.0 0.246 91.8 0.208 13.9 0.348 71.4 0.333 66.2 0.243
k-LLMmeans-FS5 82.2 0.247 92.5 0.198 13.9 0.346 71.9 0.337 65.6 0.254

### v2

# k-LLMmeans: Scalable, Stable, and Interpretable Text Clustering via LLM-based Centroids

@@ Abstract
We introduce k-LLMmeans, a novel modification of the k-means algorithm for text clustering that leverages LLM-generated summaries as cluster centroids, capturing semantic nuances often missed by purely numerical averages. This design preserves the core optimization properties of k-means while enhancing semantic interpretability and avoiding the scalability and instability issues typical of modern LLM-based clustering. Unlike existing methods, our approach does not increase LLM usage with dataset size and produces transparent intermediate outputs. We further extend it with a mini-batch variant for efficient, real-time clustering of streaming text. Extensive experiments across multiple datasets, embeddings, and LLMs show that k-LLMmeans consistently outperforms k-means and other traditional baselines and achieves results comparable to state-of-the-art LLM-based clustering—with a fraction of the LLM calls. Finally, we present a case study on sequential text streams and introduce a new benchmark dataset constructed from StackExchange to evaluate text-stream clustering methods.
@@ 1 Introduction
Text clustering is a fundamental task in natural language processing (NLP), widely applied in document organization, topic modeling, and information retrieval [37, 42]. A common approach involves generating text embeddings [13, 36, 27, 32, 8, 21] for each document, which are then clustered using traditional algorithms [33]. Among these, k-means [26] is one of the most widely used, iteratively refining cluster centroids based on the mean of assigned points. Several alternative methods exist for computing centroids [19, 7, 23], yet they often rely on purely numerical calculations. While effective, such approaches risk losing crucial contextual and semantic information, as averaging high-dimensional embeddings can obscure nuanced textual relationships [34].
@@
We propose k-LLMmeans, a modification to the k-means algorithm by leveraging large language models (LLMs) to dynamically generate cluster centroids. Instead of always computing centroids as the mean of embedding vectors (k-means), in spaced iterations we generate centroids based on textual summaries of the cluster contents. At each of these iterations, the centroid is represented as the embedding of the LLM-generated summary of the texts belonging to the cluster (summarization step). LLM-based centroids preserve key contextual and semantic aspects of the main documents in the cluster while filtering out secondary or lower-priority documents [20]. Figure 1 illustrates the k-LLMmeans algorithm with a single summarization step. The example demonstrates how this step enables the algorithm to escape local minima (common in k-means) and converge to the optimal solution.
Unlike k-means, k-LLMmeans generates a concise textual summary for every centroid at each summarization step, exposing the semantic drift of clusters in real time. This built-in transparency simplifies debugging and validation and is especially valuable in sequential or streaming scenarios, where clusters evolve over time. Figure 2 illustrates one centroid through five summarization updates. To capitalize on this property for large-scale or online data, we adapt mini-batch k-means [38] by inserting our summarization step into a mini-batch k-LLMmeans.
Several studies have explored unsupervised text clustering using LLMs [51, 15, 10, 45, 39, 43, 28], demonstrating state-of-the-art performance across various datasets and benchmarks. These approaches consistently outperform traditional clustering algorithms. Yet despite their impressive empirical gains, existing LLM-based clustering pipelines still face two practical obstacles that limit widespread adoption. First, lack of scalability. Most methods rely on fine-tuning [51], semi-supervised approaches [45], or extensive LLM queries that grow at least linearly with sample size [15, 10], quickly becoming prohibitive for real-world applications. Second, unclear optimization landscape. Most methods braid together bespoke heuristics—hard prompts, similarity thresholds, or greedy merges—without a single, well-defined loss function, leaving convergence behaviour opaque and theoretical guarantees out of reach. We address these two shortcomings with k-LLMmeans: First, by summarizing each cluster in batches we cap LLM usage independent of dataset size; and second, by inheriting the classic k-means objective we retain a provable descent property and a transparent optimization landscape. Regardless of summarization quality, the algorithm will converge to a local minimum.
@@
Our approach is not intended to outperform complex state-of-the-art LLM-based clustering methods, but rather to provide a robust, scalable and LLM-efficient enhancement to the well-established k-means algorithm, preserving its core behavior while using significantly fewer LLM calls than contemporary LLM-based clustering approaches. Fundamentally, a centroid is a numerical abstraction that represents a cluster, and we posit that an LLM-generated textual summary can serve an analogous role. Crucially, our approach preserves most of the mathematical properties of k-means, ensuring that core theoretical guarantees—such as convergence behavior, cluster compactness, and complexity— remain intact.
Figure 1: Illustration of a single-summarization k-LLMmeans algorithm. The first panel displays the text embeddings projected onto the first two principal components along with sample texts, with stars indicating initial random centroids. The second panel shows the local minimum reached by standard k-means. The third panel illustrates the summarization step, where centroids are updated based on LLM-generated summaries of the clustered texts. The final panel presents the resulting clusters after the final k-means step, showing convergence to a more optimal k-LLMmeans solution.
@@
  • In Section 3 we introduce k-LLMmeans, which enhances the classical k-means algorithm with LLM-powered centroid updates for robust text clustering; Section 4 then builds on this foundation with a mini-batch variant that brings sequential, scalable, and intrinsically interpretable clustering to streaming text data.
  • Section 5 presents extensive simulations demonstrating that k-LLMmeans consistently surpasses k-means and other traditional clustering baselines across a wide range of embeddings and LLMs, while obtaining comparable results to recent LLM-based clustering methods with substantially fewer LLM calls.
  • Section 6 shows that mini-batch k-LLMmeans decisively outperforms streaming variants of k-means, and the case study in Section 7 illustrates how its evolving centroids provide clear, human-readable insight into real-time text streams.
As a by-product of our experiments, we compile a dataset from StackExchange [41] suitable for benchmarking text-streaming analysis methods (Appendix A.2).
2 Preliminaries: k-means for text clustering
We can formalizae the text clustering problem. Given a corpus of $n$ text documents $D=\{d_{1},\cdots d_{n}\}$. Each document $d_{i}$ is represented as a $d$-dimensional embedding vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$ such that:
$\mathbf{x}_{i}=\text{Embedding}(d_{i}).$
The goal of k-means clustering is to partition these $n$ document embeddings into $k$ clusters, minimizing the intra-cluster variance. Formally, we define the clustering objective as:
$\min_{C_{1},C_{2},\dots,C_{k}}\sum_{j=1}^{k}\sum_{i\in[C_{j}]}\|\mathbf{x}_{i}-\boldsymbol{\mu}_{j}\|^{2},$ (1)
$\boldsymbol{\mu}_{j}=\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}.$ (2)
The k-means algorithm assigns each document embedding $\mathbf{x}_{i}$ to the closest centroid based on the smallest distance and updates centroids accordingly until convergence after $T$ iterations. The primary objective of the k-means algorithm is to iteratively adjust cluster centroids to minimize the within-cluster variance, effectively guiding the algorithm toward an optimal set of centroids. However, due to its sensitivity to initialization and the non-convex nature of its objective function, k-means does not guarantee convergence to the global optimum and can instead become trapped in local optima [26, 25]. Various strategies, such as k-means++ initialization and multiple restarts, have been proposed to mitigate these issues and improve the likelihood of achieving better clustering results [4].
3 k-LLMmeans
@@ 2 Preliminaries: k-means for text clustering
             // Initialize using k-means++ (omitted if already initialized)
@@
             // Update with LLM every $l$ iterations (summarization step)
@@
             // Update centroids using standard averaging (k-means step)
@@
where $p_{j}=\text{Prompt}\left(I,\{d_{z_{i}}|z_{i}\sim[C_{j}]\}_{i=1}^{m_{j}}\right)$, and $m_{j}=\min(m,|C_{j}|)$. Here, $z_{i}\sim[C_{j}]$ denotes a sampled index of the embeddings assigned to cluster $C_{j}$ (without repetitions) and $m$ is a parameter that represents the maximum number of sampled indices used to compute the cluster centroid $\boldsymbol{\mu}_{j}$. In simple terms, we update a cluster’s centroid by using the embedding of the response generated by an LLM when queried with a prompt containing a summarization instruction $I$ and a representative sample of documents from the cluster. Rather than providing all documents within the cluster as input, the LLM processes a representative sample as a context prompt. While incorporating the entire cluster is theoretically possible, it poses practical challenges due to prompt length limitations. Therefore, we propose selecting the sample cluster documents using a k-means++ sampling of the cluster embeddings. Our experiments demonstrate that this sampling process facilitates a more effective synthesis of the cluster’s content, leading to improved summaries and, consequently, more refined centroid updates. The instruction $I$ varies depending on the clustering task, but standard summarization prompts are generally sufficient. Figure 1 illustrates how k-LLMmeans augments the standard k-means algorithm by introducing few summarization steps between clustering iterations.
@@
Advantages over k-means. This novel approach can help mitigate the tendency of k-means to get stuck in local optima by dynamically adjusting centroids using semantic insights from the data. Unlike standard k-means, which relies solely on Euclidean updates, an LLM can refine centroids based on contextual meaning and high-dimensional representations, allowing for better adaptation to complex structures in text data. By periodically re-centering clusters with LLM-informed embeddings, the algorithm can escape poor local minima and achieve more coherent and semantically meaningful clusters, even when the initial k-means++ seeding is suboptimal. Apart from this modification, our algorithm adheres to the core principles of k-means, preserving its well-established properties and ensuring practical robustness. We demonstrate in Section 5.1 that k-LLMmeans consistently outperforms k-means across extensive simulations.
Advantages over advanced LLM-based clustering methods. Our approach offers three key advantages over more complex LLM-based clustering methods: (1) Optimization landscape. Grounded in the standard k-means objective, k-LLMmeans inherits provable convergence without relying on the fragile heuristics common in newer LLM-driven methods. A faulty summary merely returns the algorithm to a regular k-means local optimum, whereas competing approaches hinge on stable, format-specific LLM outputs. (2) Scalability. Unlike most state-of-the-art methods, whose LLM usage complexity grows with sample size [15, 10] or require fine-tuning [51], our algorithm only performs $O(kl)$ LLM calls, with even small $l$ values yielding substancial performance gains. (See Section 5.1). (3) Interpretability. By replacing numerical centroids with LLM-generated textual summaries, k-LLMmeans turns each centroid into a concise, human-readable synopsis of its cluster. Practitioners can therefore monitor how cluster semantics evolve across iterations without any post-hoc labeling or visualization. While comparable insights can be extracted indirectly from other clustering algorithms, k-LLMmeans embeds this interpretability natively. Figure 2 illustrates how these summaries progressively sharpen as the algorithm converges.
@@ 4 Mini-batch k-LLMmeans
input : $\ \{D_{1},\cdots,D_{b}\},k,I,m,l,T$ // $b$ batches of documents
@@
       // Update centroids proportional to cluster and batch sizes
@@
Mini-batch k-means [38] is an efficient strategy for large-scale text clustering that processes small, randomly sampled mini-batches instead of the full dataset. This approach substantially reduces memory usage and computational cost, making it well suited for continuously generated text streams—such as those from social media, news, or customer feedback—where data must be clustered incrementally without full dataset access. Mini-batch k-means exhibits convergence properties comparable to standard k-means while offering superior scalability.
@@
Although numerous streaming clustering methods that do not rely on LLMs have been studied [40, 2, 35, 3, 1, 31], only a few have incorporated LLMs [43, 28]. Moreover, existing offline LLM-based clustering approaches face scalability issues, highlighting the need for scalable LLM-driven clustering in an online setting. To address this gap, we introduce mini-batch k-LLMmeans, which directly extends mini-batch k-means by incorporating minimal LLM usage during centroid updates. Algorithm 2 details how mini-batch k-LLMmeans sequentially receives $b$ batches of documents $D_{1},\dots D_{b}$ where each batch contains a set of documents (these batches can either be random samples from a large corpus or represent sequential data). It processes each batch sequentially with k-LLMmeans and updates centroids incrementally using a weighted rule like mini-batch k-means. Our mini-batch k-LLMmeans algorithm preserves the desirable properties of mini-batch k-means, with low memory and LLM usage. Section  6 shows that it also outperforms in simulations.
5 Static experiments
We use four benchmark datasets that span diverse domains and classification granularities: Bank77 [9], CLINC [24], GoEmo [12] and MASSIVE (domain and intent) [16]. See Appendix A.1 for detailed descriptions of the datasets. We evaluate our k-LLMmeans algorithm on each of the four static (non-streaming) dataset using the known number of clusters and performing 120 centroid-update iterations. To demonstrate the robustness of our approach, we compute embeddings with four different pretrained models: DistilBERT [36], e5-large [46], S-BERT [34], and text-embedding-3-small [30]. For the LLM component, we use GPT-3.5-turbo, GPT-4o, Llama-3.3, Claude-3.7 and DeepSeek-V3.
For the instruction task $I$, we employ a simple summarization prompt that adapted to each dataset. For example, for Bank77 we use the prompt: “The following is a cluster of online banking questions. Write a single question that represents the cluster concisely.” We evaluate four variants of k-LLMmeans, differing in the number of summarization steps and the prompt configuration. The single variant uses a single summarization step with all cluster documents as input ($l=60$), while the multiple variant performs five summarization steps with shorter prompts ($l=20$), again using all cluster documents each time. Additionally, we explore few-shot variants (FS-single and FS-multiple), where each prompt includes $m=10$ randomly selected documents instead of the full cluster.
As traditional clustering baselines, we consider k-means, k-medoids [23], spectral clustering [29], agglomerative clustering [22], and Gaussian Mixture Models (GMM) [11]. We use ground-truth number of clusters and ten different seeds. For advanced LLM-based methods, we compare with results obtained by Feng et al. [15] on ClusterLLM [51], IDAS [10], and LLMEdgeRefine [15].
5.1 Results with static data
Table 1: Average Accuracy (ACC) and Normalized Mutual Information (NMI) for four k-LLMmeans variants using GPT-4o, compared against traditional clustering baselines, with text-embedding-3-small embeddings, evaluated on four benchmark datasets.
Dataset/Method BANK77 CLINC GoEmo Massive (D) Massive (I)
ACC NMI ACC NMI ACC NMI ACC NMI ACC NMI
k-LLMmeans
   single 67.1 83.6 78.1 92.5 24.0 22.3 61.1 69.6 54.0 73.0
   multiple 66.7 83.6 79.1 92.8 24.1 22.1 60.6 69.5 55.8 73.5
   FS-single 67.5 83.8 79.2 92.8 23.0 21.9 60.8 69.4 55.3 73.6
   FS-multiple 67.9 84.1 80.2 93.1 24.2 22.3 63.2 70.6 56.4 73.9
k-means 66.2 83.0 77.3 92.0 20.7 20.5 59.4 67.9 52.9 72.4
k-medoids 41.7 69.5 49.3 77.7 15.7 15.9 37.6 38.8 35.9 52.4
GMM 67.7 83.0 78.9 92.7 21.5 20.6 56.2 68.6 53.9 73.2
Agglomerative 69.9 83.7 81.0 92.5 15.8 14.1 62.8 67.1 56.6 70.5
Spectral 68.2 83.3 76.3 90.9 17.6 15.2 61.5 67.0 56.7 71.4